OpenStudy (anonymous):

Two gears are connected and are rotating simultaneously. The smaller gear has a radius of 4 inches, and the larger gear has a radius of 7 inches. two circles touching at one point. Larger circle has radius of 7 inches. Smaller circle has radius of 4 inches. Part 1: What is the angle measure, in degrees and rounded to the nearest tenth, through which the larger gear has rotated when the smaller gear has made one complete rotation? Part 2: How many rotations will the smaller gear make during one complete rotation of the larger gear?

OpenStudy (anonymous):

@phi @myininaya @ganeshie8 @mathmath333

OpenStudy (anonymous):

can someone help pls

OpenStudy (phi):

I would find the circumference of each circle as a first step.

OpenStudy (anonymous):

ok

OpenStudy (mathmath333):

let r1 be radius for small circle and let r2 be radius for small circle therefore length of arc S \(\large\tt \begin{align} \color{black}{s_1=r_1\theta_1\\~\\ s_2=r_2\theta_2\\~\\ }\end{align}\) u have find \(\large\tt \begin{align} \color{black}{\theta_2 }\end{align}\)when \(\large\tt \begin{align} \color{black}{\theta_1 =2\pi\\~\\ }\end{align}\) \(\large\tt \begin{align} \color{black}{s_1=r_1\theta_1\\~\\ s_1=4(2\pi)\\~\\ }\end{align}\) u have to find the angle through which the larger gear has rotated when the smaller gear has made one complete rotation. \(\large\tt \begin{align} \color{black}{thus~~s_1=s_2\\~\\ s_2=r_2\theta_2\\ 8\pi=7\theta_2\\ \theta_2=\dfrac{8\times 180}{7}\\ =205~~ degrees~~ }\end{align}\)

OpenStudy (anonymous):

ok thx